OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (2, –5) and (–4, 1)?

OpenStudy (anonymous):

If two lines are parallel that means that they have the same slope. The slope of the line crossing these points is: m = (y2 - y1)/(x2 - x1) =(1 - (-5))/(-4 - 2) =6/(-6) =-1 So any line that has a slope of -1 is parallel. Is this a multiple choice?